OpenStudy (anonymous):

(Alg 2) The pH of a liquid is a measure of how acidic or basic it is. The concentration of hydrogen ions in a liquid is labeled [H^+]. Use the formula pH = -log [H^+] to find the pH level, to the nearest tenth, of a liquid with [H^+] about 6.5 x 10^-3.

OpenStudy (agent0smith):

\[\large pH = -\log [H^+] \] and \[\large [H^+] = 6.5 \times 10^{-3}\] so \[\large pH = -\log (6.5) (10^{-3})\] Now try using the log law \[\log(a)(b) = \log a+\log b\]

OpenStudy (agent0smith):

(or you can do it on a calculator, as you'll need to use it to find the pH anyway)

OpenStudy (anonymous):

@agent0smith I got 8.1 x 10^-4?

OpenStudy (agent0smith):

Show me how you got that...

OpenStudy (anonymous):

I inputted -log(6.5)(10^-3) on my graphing calculator and then divided by -1. My teacher showed us how to do a different kind of equation so I'm super confused with this problem! She was like having us divide by -1 first and then rewrite into exponents, then in scientific notation but I'm not sure if the question wants us to do that. @agent0smith

OpenStudy (agent0smith):

If you do it on a calculator, make sure you use parentheses! Do it as \[\LARGE -1 \times \log( 6.5 \times 10^{-3 } )\]

OpenStudy (anonymous):

Ok! Just tried it. I got -8.1 x 10^-4? @agent0smith

OpenStudy (agent0smith):

You must not be entering it into your calculator correctly... it should be 2.18.

OpenStudy (anonymous):

Oh, I don't think the -1 thing worked for my calculator. I just looked at the wolfram alpha thing and I just got the 2.18 by just doing the -log. I wonder why that is haha! But thank you for your help! I appreciate it!

OpenStudy (agent0smith):

Don't enter -log(6.5)(10^-3), enter -log(6.5x10^-3)

OpenStudy (anonymous):

Oh that's what happened! ^^ Thank you! :)

OpenStudy (agent0smith):

No prob :) I see how you got your 8.1 answer, if you enter -log(6.5)(10^-3), then your calculator does -log(6.5) and multiplies it by (10^-3), it doesnt find -log(6.5x10^-3).

OpenStudy (anonymous):

That makes sense! I will remember that for my test. Thanks again :) I've been stressing over that problem haha! @agent0smith

OpenStudy (agent0smith):

No prob :)